at present father's age is thrice that of his son 6 years back his age was four times than that of his son .what will be the ratio of their ages after 6 years ?
Answer by ankor@dixie-net.com(22740)

let f = father's present age
let s = son's
:
Write an equation for each statement
:
At present father's age is thrice that of his son.
f = 3s
6 years back his age was four times than that of his son.
f - 6 = 4(s-6)
f - 6 = 4s - 24
f = 4s - 24 + 6
f = 4s - 18
Replace f with 3s from the 1st statement
3s = 4s - 18
18 = 4s - 3s
s = 18 yrs is son's age
then
3(18) = 54 yrs is father's age
:
What will be the ratio of their ages after 6 years?
then dad will be 60 and son will 24
60:24 reduced to 5:2 ratio